### v1

# Comparison of Extended Lubrication Theories for Stokes Flow

Abstract
Lubrication theory leverages the assumption of a long and thin fluid domain to formulate a linearized approximation to the Navier-Stokes equations. Models in extended lubrication theory consider relaxations of the thin film assumption, leading to the inclusion of higher order terms. However, such models are sensitive to large surface gradients which lead the assumptions of the model to break down. In this paper, we present a formulation of extended lubrication theory, and compare our models with several existing models, along with the numerical solution to the Stokes equations. The error in pressure and velocity is characterized for a variety of fluid domain geometries. Our results indicate that the new solution is suitable for a wide range of geometries. Nonetheless, the magnitude of surface variation and the length scale ratio are both important factors influencing the accuracy of the extended lubrication theory models.
1 Introduction
The classical lubrication assumption of a long and thin fluid domain is used to justify neglecting the inertial velocity and cross-film pressure gradient terms in the Navier-Stokes equations. The resulting Reynolds equation is a dramatic simplification of the original governing equations. More recently, variations on the thin film assumption have given rise to higher order approximations for lubrication theory [16, 15, 14]. These extended lubrication theory models retain certain linearity properties as in classical lubrication theory, while nevertheless being applicable to a broader scope of geometries in the low Reynolds number limit.
Previous research on the flows modeled by lubrication theory has highlighted the sensitivity of lubrication theory to variations of the fluid film height. In particular, in the limit of zero Reynolds number flows, discrepancies between lubrication theory and Stokes flow increase with both the magnitude and frequency of large gradients in the height [4, 2, 5]. The total pressure drop modeled by the Reynolds equation may be underestimated due to additional pressure losses at points of sudden expansion that are not accurately modeled [6, 5]. Furthermore, notable flow features of corner recirculation seen in Stokes flow are not accurately captured with the Reynolds equation [1, 2, 13].
In this paper, we consider classical lubrication theory (the Reynolds equation), and more recent variations on lubrication theory which we refer to as extended lubrication theory [14] and/or perturbed lubrication theory [16, 10, 15] in the case of models based on perturbation theory. We asses the performance of these models at low Reynolds numbers by comparison to the Stokes equations. We propose a formulation of extended lubrication theory that builds on the model of Takeuchi & Gu (2019); our approach modifies the treatment of the velocity boundary conditions and ensures incompressibility. The pressure and velocity obtained from each model are compared in a variety of geometries: a logistic step, a backward facing step, and two versions of a triangular slider that we refer to as positive and negative based on the texture to be defined later on. Through these examples, we assess the sensitivity of each model to varying magnitudes of surface variation and discontinuities in the surface gradient.
Our results indicate that the extended and perturbed lubrication theory models provide significant improvements on classical lubrication theory in the case of small to moderate surface variation. For the logistic step, our proposed solution in extended lubrication theory has the smallest error in velocity for small slopes, and the smallest error in pressure at moderate slopes. For the triangular slider, the perturbed solutions have significantly reduced error in pressure, particularly in the case of negative texturing. In general, we find that for examples with a large surface gradient, the extended and perturbed solutions may over-correct the cross-film pressure gradient.
2 Background
2.1 Lubrication Theory
We begin with a review of classical lubrication theory. This treatment may be found in various textbooks [8], and we include it here to establish notation and to make the presentation self-contained. The Navier-Stokes equations,
(1) $\frac{\partial p}{\partial x}$ $=\nu\Big(\frac{\partial^{2}u}{\partial x^{2}}+\frac{\partial^{2}u}{\partial y^{2}}\Big)-\Big(u\frac{\partial u}{\partial x}+v\frac{\partial u}{\partial y}\Big),$
(2) $\frac{\partial p}{\partial y}$ $=\nu\Big(\frac{\partial^{2}v}{\partial x^{2}}+\frac{\partial^{2}v}{\partial y^{2}}\Big)-\Big(u\frac{\partial v}{\partial x}+v\frac{\partial v}{\partial y}\Big),$
subject to incompressibility,
(3) $\frac{\partial u}{\partial x}+\frac{\partial v}{\partial y}=0,$
serve as the governing equations for a two dimensional incompressible fluid. The pressure is $p$, the velocity is $(u,v)$, and the kinematic viscosity is $\nu=\eta/\rho$, where $\eta$ is the bulk viscosity and $\rho$ is the constant density.
Lubrication theory provides a linearized approximation to the Navier-Stokes equations valid for geometries in which the fluid is constrained to a relatively thin domain. For the two dimensional fluid domain $(x,y)\in[x_{0},x_{f}]\times[0,h(x)]$, the characteristic length scales are $L_{x}=x_{f}-x_{0}$ and $L_{y}=\max h(x)>0$. Given a prescribed constant flux $\mathcal{Q}\neq 0$, the characteristic velocities are $U_{*}=\mathcal{Q}/L_{y}$ and $V_{*}=\mathcal{Q}/L_{x}$, and the characteristic pressure is $P_{*}=\eta U_{*}L_{x}/L_{y}^{2}$. The nondimensionalized variables are $\bar{x}=x/L_{x}$, $\bar{y}=y/L_{y}$, $\bar{u}=u/U_{*}$, $\bar{v}=v/V_{*}$, and $\bar{p}=p/P_{*}$. The Reynolds number is $\text{Re}=\rho U_{*}L_{y}/\eta$ and the length scale ratio is denoted $\varepsilon=L_{y}/L_{x}$. The Navier-Stokes equations Eqs. 1 and 2 in dimensionless terms are then,
(4) $\frac{\partial\bar{p}}{\partial\bar{x}}$ $=\Big(\varepsilon^{2}\frac{\partial^{2}\bar{u}}{\partial\bar{x}^{2}}+\frac{\partial^{2}\bar{u}}{\partial\bar{y}^{2}}\Big)-\varepsilon^{2}\text{Re}\Big(\bar{u}\frac{\partial\bar{u}}{\partial\bar{x}}+\bar{v}\frac{\partial\bar{u}}{\partial\bar{y}}\Big),$
(5) $\frac{\partial\bar{p}}{\partial\bar{y}}$ $=\varepsilon^{2}\Big(\varepsilon^{2}\frac{\partial^{2}\bar{v}}{\partial\bar{x}^{2}}+\frac{\partial^{2}\bar{v}}{\partial\bar{y}^{2}}\Big)-\varepsilon^{4}\text{Re}\Big(\bar{u}\frac{\partial\bar{v}}{\partial\bar{x}}+\bar{v}\frac{\partial\bar{v}}{\partial\bar{y}}\Big),$
with incompressibility Eq. 3 which is unchanged.
The lubrication assumptions $\varepsilon\ll 1$ and $\varepsilon^{2}\text{Re}\ll 1$ characterize a long and thin fluid with a small scaled Reynolds number. These assumptions yield an approximation to Eqs. 4 and 5 with the resulting equations,
(6) $\frac{\partial p}{\partial x}$ $=\eta\frac{\partial^{2}u}{\partial y^{2}},$
(7) $\frac{\partial p}{\partial y}$ $=0.$
Together with incompressibility Eq. 3, Eqs. 6 and 7 constitute the governing equations for lubrication theory.
We assume the no-slip boundary condition at the fluid-surface interfaces $y=0$ and $y=h(x)$. Without loss of generality, the velocity boundary conditions are,
(8) $u(x,0)=\mathcal{U}$ $u(x,h(x))=0,$
(9) $v(x,0)=0$ $v(x,h(x))=0,$
where $\mathcal{U}$ is the constant relative velocity between the surfaces. We prescribe a constant flux $\mathcal{Q}$, corresponding to the inlet pressure gradient, and set the outlet pressure to be zero,
(10) $\frac{dp}{dx}\bigg|_{x_{0}}=\frac{-12\eta\mathcal{Q}}{\big[h(x_{0})\big]^{3}}+\frac{6\eta\mathcal{U}}{\big[h(x_{0})\big]^{2}},$ $p(x_{L},y)=0.$
The velocity $u=u_{\text{Re}}$ is determined through integration of Eq. 6 and applying the boundary conditions Eq. 8,
(11) $u(x,y)=\frac{1}{2\eta}\frac{dp}{dx}\Big(y^{2}-h(x)y\Big)+\frac{\mathcal{U}}{h(x)}\Big(h(x)-y\Big).$
The velocity $v=v_{\text{Re}}$ is determined from incompressibility Eq. 3 and applying the boundary conditions Eq. 9,
(12) $v(x,y)=\frac{-1}{6\eta}\frac{d^{2}p}{dx^{2}}y^{3}+\frac{1}{2}\Bigg(\frac{1}{2\eta}\Big(\frac{d^{2}p}{\partial x^{2}}h(x)+\frac{dp}{dx}\frac{dh}{dx}\Big)-\frac{\mathcal{U}}{[h(x)]^{2}}\frac{dh}{dx}\Bigg)y^{2}.$
The condition of constant flux $\mathcal{Q}$,
(13) $\mathcal{Q}=\int_{0}^{h(x)}u(x,y)\,dy=\frac{-1}{12\eta}\Bigg(\Big[h(x)\Big]^{3}\frac{dp}{dx}-6\eta\mathcal{U}h(x)\Bigg),$
and the boundary condition $v(x,h)=0$ are satisfied exactly when $p(x)$ satisfies the Reynolds equation,
(14) $\frac{d}{dx}\Bigg[\Big[h(x)\Big]^{3}\frac{dp}{dx}\Bigg]=6\eta\,\mathcal{U}\frac{dh}{dx}.$
The solution of pressure $p=p_{\text{Re}}$ for Reynolds equation is determined using a second-order accurate finite difference scheme, see Appendix A.
2.2 Perturbed Lubrication Theory
The dimensionless Navier-Stokes equations Eqs. 4 and 5 motivate the approach of an $\varepsilon$-perturbed solution. Given a sufficiently small Reynolds number such that $\varepsilon^{2}\text{Re}\ll 1$, perturbed lubrication theory assumes a solution of pressure and velocity of the form,
(15) $\bar{p}(\bar{x},\bar{y})$ $=P_{0}+\varepsilon^{2}P_{2}+\varepsilon^{4}P_{4},$
(16) $\bar{u}(\bar{x},\bar{y})$ $=U_{0}+\varepsilon^{2}U_{2}+\varepsilon^{4}U_{4},$
(17) $\bar{v}(\bar{x},\bar{y})$ $=V_{0}+\varepsilon^{2}V_{2}+\varepsilon^{4}V_{4}.$
The order $\varepsilon^{k}$ solution ($\varepsilon^{k}$-PLT) assumes a long and thin domain characterized by $\varepsilon^{k}\sim\mathcal{O}(1)$ and such that higher powers of $\varepsilon$ are negligible; the order $\varepsilon^{0}$-PLT solution is the solution of lubrication theory and the Reynolds equation. This approach of perturbed solutions has been discussed e.g. by Venner et. al. [16], Marusic et al. [10] and Tavakol et al. [15].
On substituting the ansatz Eqs. 15, 16 and 17 into Eqs. 4 and 5 at order $\varepsilon^{2}$, the velocity and pressure satisfy the following reduced equations,
(18) $\frac{\partial P_{2}}{\partial\bar{x}}$ $=\frac{\partial^{2}U_{0}}{\partial\bar{x}^{2}}+\frac{\partial^{2}U_{2}}{\partial\bar{y}^{2}},$
(19) $\frac{\partial P_{2}}{\partial\bar{y}}$ $=\frac{\partial^{2}V_{0}}{\partial\bar{y}^{2}}.$
Likewise at order $\varepsilon^{4}$,
(20) $\frac{\partial P_{4}}{\partial\bar{x}}$ $=\frac{\partial^{2}U_{2}}{\partial\bar{x}^{2}}+\frac{\partial^{2}U_{4}}{\partial\bar{y}^{2}},$
(21) $\frac{\partial P_{4}}{\partial\bar{y}}$ $=\frac{\partial^{2}V_{2}}{\partial\bar{y}^{2}}+\frac{\partial^{2}V_{0}}{\partial\bar{x}^{2}}.$
The velocity fields $(U_{k},V_{k})$ additionally all satisfy incompressibility Eq. 3.
We assume the boundary conditions Eqs. 8, 9 and 10. Then $U_{0}$ and $V_{0}$ are the dimensionless forms of Eqs. 11 and 12. and $P_{0}$ is the dimensionless form of $p_{\text{Re}}$ obtained numerically. We define the flux $\mathcal{Q}$ of the $\varepsilon^{k}$-PLT solution to be consistent with that of the order $\varepsilon^{0}$ solution. Then in general, the pressure drop $\Delta\mathcal{P}$ of the $\varepsilon^{k}$-PLT solution will deviate from that of the order $\varepsilon^{0}$ solution.
The method of perturbed solutions is appealing in that we obtain exact expressions for the velocities $U_{k}$ and $V_{k}$ involving only $y$, the height function $h(x)$ and its derivatives. The expressions for pressure $P_{k}$ do involve integrals requiring numerical integration for an arbitrary height function. The complete expressions for $U_{k}$, $V_{k}$ and $P_{k}$ at orders $\varepsilon^{2}$ and $\varepsilon^{4}$ are presented in the appendix Appendix C. Similar expressions up to various orders appear in [16, 10, 15]. Here we have provided corrected expressions for $V_{2}$ and $U_{4}$ in [15], and we include an expression for $V_{4}$ which is not otherwise explicitly written in [16, 10, 15].
2.3 Extended Lubrication Theory
An alternative extended lubrication theory model (T.G.-ELT) is proposed by Takeuchi and Gu [14], which retains terms of order up to $\varepsilon^{2}$ in Eqs. 4 and 5 to result in the governing equations,
(22) $\frac{\partial p}{\partial x}$ $=\eta\frac{\partial^{2}u}{\partial y^{2}},$
(23) $\frac{\partial p}{\partial y}$ $=\eta\frac{\partial^{2}v}{\partial y^{2}},$
with incompressibility Eq. 3 and the boundary conditions Eqs. 10, 8 and 9. Note that the second-order term $\varepsilon^{2}\frac{\partial^{2}\bar{u}}{\partial\bar{x}^{2}}$ appearing in Eq. 4 is not included in the T.G.-ELT formulation, since they argue this may be neglected in the case of small surface gradients. As with the perturbed solutions, the flux $\mathcal{Q}$ is assumed to be equal the Reynolds flux, and in general, the pressure drop $\Delta\mathcal{P}$ may deviate from that of the solution to the Reynolds equation.
According to the T.G.-ELT assumptions, the pressure gradient $\frac{\partial^{2}p}{\partial y\partial x}$ is negligible provided that the condition $\varepsilon^{2}\ll 1$ is satisfied [14]. Since the $x$ momentum equations of lubrication theory and T.G.-ELT are the same, integration in $y$ of Eq. 22 gives $u=u_{\text{Re}}$ as in Eq. 11, and applying incompressibility gives $v=v_{\text{Re}}$ as in Eq. 12. However, the $y$ momentum equations are different; integration in $y$ of Eq. 23 gives the pressure decomposition,
(24) $p(x,y)=p_{\text{Re}}(x)+p_{\text{Adj}}(x,y)+\eta\varsigma(x),$
where $p_{\text{Adj}}(x,y)=\eta\frac{\partial v_{\text{Re}}}{\partial y}$ and $\varsigma(x)$ is an integral constant. That is,
(25) $p_{\text{Adj}}(x,y)=\frac{-1}{2}\frac{d^{2}p_{\text{Re}}}{dx^{2}}y^{2}+\Bigg(\frac{1}{2}\bigg(\frac{d^{2}p_{\text{Re}}}{dx^{2}}h(x)+\frac{dp_{\text{Re}}}{dx}\frac{dh}{dx}\bigg)-\frac{\mathcal{U\eta}}{\big[h(x)\big]^{2}}\frac{dh}{dx}\Bigg)y.$
3 Velocity-Adjusted Extended Lubrication Theory
In [14], the integral constant $\varsigma(x)$ is assumed negligible on account of $\varepsilon^{2}\ll 1$. In the construction of the velocity to follow, we remove this assumption and allow $\varsigma(x)\neq 0$ to ensure the velocity derived from $p(x,y)$ satisfies the boundary conditions and incompressibility, even in regions where the T.G.-ELT conditions are violated. The inclusion of the term $\varsigma(x)$ leads to $p$, $u$, and $v$ each having different expressions compared with the T.G.-ELT solution; we refer to the new solution with $\varsigma(x)\neq 0$ as the velocity adjusted extended lubrication theory solution (VA-ELT).
Since $p_{\text{Adj}}(x,y)$ is polynomial in $y$, the $x$ momentum equation Eq. 22 may be integrated exactly in $y$ to obtain $u$. We utilize the integral constant $\varsigma(x)$ to impose the boundary conditions and incompressibility independent of the local geometry. Applying the boundary conditions Eq. 8 for $u$,
(26) $u(x,y)=u_{\text{Re}}(x,y)-\frac{1}{24\eta}\frac{d^{3}p_{\text{Re}}}{dx^{3}}(y^{4}-h^{3}y)+\frac{1}{12\eta}\frac{d^{2}}{dx^{2}}\Bigg[h\frac{dp_{\text{Re}}}{dx}\Bigg](y^{3}-h^{2}y)\\ +\frac{\mathcal{U}}{6}\frac{d^{2}}{dx^{2}}\Bigg[\frac{1}{h}\Bigg](y^{3}-h^{2}y)+\frac{1}{2}\frac{d\varsigma}{dx}(y^{2}-hy).$
Then applying incompressibility and integrating in $y$ with the boundary conditions Eq. 9 for $v$,
(27) $v(x,y)=v_{\text{Re}}(x,y)+\frac{1}{24\eta}\Bigg(\frac{d^{4}p_{\text{Re}}}{dx^{4}}\bigg(\frac{y^{5}}{5}-\frac{h^{3}y^{2}}{2}\bigg)-\frac{3h^{2}y^{2}}{2}\frac{d^{3}p_{\text{Re}}}{dx^{3}}\frac{dh}{dx}\Bigg)\\ -\frac{1}{12\eta}\Bigg(\frac{d^{3}}{dx^{3}}\bigg[h\frac{dp}{dx}\bigg]\bigg(\frac{y^{4}}{4}-\frac{h^{2}y^{2}}{2}\bigg)-\frac{d^{2}}{dx^{2}}\bigg[h\frac{dp_{\text{Re}}}{dx}\bigg]\frac{dh}{dx}hy^{2}\Bigg)\\ +\frac{\mathcal{U}}{6}\Bigg(\frac{y^{4}-2h^{2}y^{2}}{4h^{2}}\frac{d^{3}h}{dx^{3}}+\frac{3y^{4}-2h^{2}y^{2}}{2h^{4}}\bigg[\frac{dh}{dx}\bigg]^{3}+\frac{4h^{2}y^{2}-3y^{4}}{2h^{3}}\frac{dh}{dx}\frac{d^{2}h}{dx^{2}}\Bigg)\\ -\frac{d^{2}\varsigma}{dx^{2}}\bigg(\frac{y^{3}}{6}-\frac{hy^{2}}{4}\bigg)+\frac{d\varsigma}{dx}\frac{dh}{dx}\frac{y^{2}}{4}.$
Observe that if we were to assume $\varsigma(x)=0$, the velocity Eqs. 26 and 27 is not necessarily incompressible and Eq. 27 may not satisfy $v(x,h)=0$.
The integral constant $\varsigma(x)$ is determined by asserting a constant flux. We prescribe the flux $\mathcal{Q}=\mathcal{Q}_{\text{Re}}$ as in Eq. 13. Then $\varsigma(x)$ satisfies,
(28) $\frac{d\varsigma}{dx}=\frac{3h^{2}}{20\eta}\frac{d^{3}p_{\text{Re}}}{dx^{3}}-\frac{h}{4\eta}\bigg(h\frac{d^{3}p_{\text{Re}}}{dx^{3}}+2\frac{d^{2}p_{\text{Re}}}{dx^{2}}\frac{dh}{dx}+\frac{dp_{\text{Re}}}{dx}\frac{d^{2}h}{dx^{2}}\bigg)+\frac{\mathcal{U}}{2}\bigg(\frac{1}{h}\frac{d^{2}h}{dx^{2}}-\frac{2}{h^{2}}\bigg[\frac{dh}{dx}\bigg]^{2}\Bigg).$
For an arbitrary height function, numerical integration is required to obtain $\varsigma(x)$. The boundary condition $\varsigma(x_{L})=0$ corresponds to $p(x_{L},y)=p_{\text{Re}}(x_{L},y)=0$. The total pressure drop is then $\Delta\mathcal{P}=\Delta\mathcal{P}_{\text{Re}}+\varsigma(x_{0})$.
4 Results
We now compare the lubrication models for various geometries and examine the solutions of velocity and pressure for a variety of textured slider examples. The Reynolds, T.G.-ELT, VA-ELT, $\varepsilon^{2}$-PLT and $\varepsilon^{4}$-PLT solutions are compared against the Stokes solution Appendix B. The accuracy of pressure and velocity solutions are characterized using the relative percent error in $L_{2}$ norm. For an appropriate comparison of solutions from the various models, the flux $\mathcal{Q}$ is kept constant and the pressure drops $\Delta\mathcal{P}$ are presumed to vary.
Note, for the T.G.-ELT method, the prescribed flux $\mathcal{Q}$ corresponds to the pressure gradient $\frac{\partial p}{\partial x}$ assigned at the inlet. However – subject to the magnitude of surface variation – this flux may not be conserved throughout the domain because incompressibility is not strictly enforced.
4.1 The logistic step
First, we consider the logistic step, shown in Fig. 1. The height function is given by,
(29) $h(x)=H_{\text{in}}+\frac{H_{\text{out}}-H_{\text{in}}}{1+e^{\lambda\big(L/2-x\big)}},$
where $H_{\text{in}}\geq H_{\text{out}}>0$ and $\lambda>0$. The logistic step allows us to vary the magnitude of surface gradient while keeping a fixed $\varepsilon=L_{y}/L_{x}$ and step expansion ratio $H_{\text{in}}/H_{\text{out}}$. We consider solutions for $2\leq\lambda\leq 32$, and with fixed $H_{\text{in}}=2$, $H_{\text{out}}=1$, $L=16$, $\nu=1$, $\mathcal{U}=0$ and $\mathcal{Q}=1$. Then the length scale ratio is $\varepsilon\simeq 1/8$, and the maximum magnitude surface gradient is $\big|\frac{dh}{dx}\big|=\lambda/4$.
Figure 1: Schematic of the logistic step.
The pressure and velocity for the logistic step with $\lambda=8$ are shown in Fig. 2. The pressure contours for the Stokes solution exhibit significant $\frac{\partial p}{\partial y}$ in a wider region surrounding the large surface gradient compared with the ELT and $\varepsilon^{k}$-PLT solutions, whereas the ELT and $\varepsilon^{k}$-PLT solutions overestimate $\frac{\partial p}{\partial y}$ in the localized region of surface variation. This leads to flow recirculation being observed at smaller $\lambda$ in the ELT and $\varepsilon^{k}$-PLT solutions than with the Stokes solution.
(a) Pressure contours $p(x,y)$
(b) Velocity streamlines $(u,v)$
Figure 2: Pressure and velocity solutions for the logistic step with $\lambda=8$. The ELT and $\varepsilon^{k}$-PLT solutions overestimate $\frac{\partial p}{\partial y}$ in the vicinity of the surface variation, leading to flow recirculation which is not observed in the Stokes solution at this moderate $\lambda$.
The errors in velocity and pressure compared with the Stokes solution are shown in Fig. 3. For the wide range of $\lambda<32$, the VA-ELT solution of pressure is significantly more accurate than the Reynolds and T.G.-ELT solutions (Fig. 3). The T.G.-ELT solution of pressure yields a marginal improvement over the Reynolds pressure. For $2<\lambda<8$, the $\varepsilon^{2}$-PLT solution is best at approximating the pressure, and the VA-ELT solution is best at approximating the velocity. The $\varepsilon^{2}$-PLT solution performs better than the $\varepsilon^{4}$-PLT solution for all $\lambda>2$. For $\lambda\geq 32$, the Reynolds solution has the smallest error in velocity and pressure; the Reynolds solution has the most stable error across the range of $\lambda$. Overall, both the $\varepsilon^{2}$-PLT and VA-ELT solutions perform well for the logistic step up to moderate surface variation. We note however, that all the lubrication models discussed here break down in the limit $\lambda\to\infty$ of a discontinuous step, see Appendix D.
Figure 3: Relative percent error in velocity $(u,v)$ and pressure $p(x,y)$ compared with the Stokes solution for the logistic step at varying slopes, $\lambda/4=\max|\frac{dh}{dx}|$. For $2<\lambda\leq 6$, the VA-ELT solution has the smallest error in velocity and the $\varepsilon^{2}$-PLT solution has the smallest error in pressure.
4.2 The triangular slider
Next, we consider a triangular slider, shown in Fig. 4. The height function is given by,
(30) $h(x)=\begin{cases}H_{\text{in}}&x\leq L_{\text{in}}\\ H-\frac{H-H_{\text{in}}}{L_{\text{a}}}\big(L_{\text{in}}+L_{a}-x\big)&L_{\text{in}}\leq x\leq L_{\text{in}}+L_{a}\\ H_{\text{out}}-\frac{H_{\text{out}}-H}{L_{b}}\big(L-L_{\text{out}}-x\big)&L_{\text{in}}+L_{a}\leq x\leq L-L_{\text{out}}\\ H_{\text{out}}&L-L_{\text{out}}\leq x\leq L.\end{cases}$
We consider varying central vertex heights $\frac{1}{16}\leq H\leq 2$, and with fixed $L_{a}=1.25$, $L_{b}=0.75$, $L_{\text{in}}=L_{\text{out}}=7$, $H_{\text{in}}=H_{\text{out}}=1$, $\nu=1$, $\mathcal{U}=0$ and $\mathcal{Q}=1$. The maximum magnitude surface gradient is $|\frac{dh}{dx}|=|1-H|/0.75$, and the length scale ratio is $\varepsilon\simeq\max(1,H)/16$. This configuration is similarly considered in [15] for the case of negative texturing $(H<1)$ with the $\varepsilon^{k}$-PLT solutions.
Figure 4: Schematic of the triangular slider
The pressure and velocity solutions to the triangular slider with $H=0.5$ (following [15]) and $H=2$ are shown in Figs. 5 and 6 respectively. This class of examples demonstrates the limitations of the lubrication models in the case of a discontinuous surface gradient. At the discontinuities in $\frac{dh}{dx}$, the pressure gradient $\frac{d^{2}p_{\text{Re}}}{dx^{2}}$ and velocity $v_{\text{Re}}$ of the Reynolds solution are not necessarily continuous. Consequently, the pressure and velocity solutions from ELT and $\varepsilon^{k}$-PLT may not be continuous at the discontinuities of $\frac{dh}{dx}$. Note that $u_{\text{Re}}$ and $\frac{dp_{\text{Re}}}{dx}$ are continuous for the triangular slider because $h(x)$ is continuous; owing to the lubrication assumptions, $v_{\text{Re}}$ is typically small enough in magnitude compared to $u_{\text{Re}}$ that the discontinuity in $v_{\text{Re}}$ is not significant. However, for larger $H$, the discontinuity in $v_{\text{Re}}$ is more pronounced and leads to a noticeable loss of smoothness in the ELT and $\varepsilon^{k}$-PLT solutions.
In the case of a sufficiently small interior angle at the triangle apex, the Stokes solution is known to exhibit a sequence of eddies receding into the corner [11, 3]. For $H=2$ as in Fig. 6, the first eddy is clearly observed in the Stokes solution. For the ELT and $\varepsilon^{k}$-PLT solutions, flow recirculation may occur if $H$ and $\frac{dh}{dx}$ are sufficiently large. However, because the velocities for these models are not continuous at the discontinuities of $\frac{dh}{dx}$, corner eddies do not occur as in the Stokes solution. In fact, as seen in Fig. 6 with the $\varepsilon^{k}$-PLT solutions, the velocity in the steeper half of the triangle exhibits flow recirculation at smaller $H$ than the shallower half. That is, if $H$ is large enough and $\frac{dh}{dx}$ is discontinuous in magnitude, the Reynolds, ELT and $\varepsilon^{k}$-PLT solutions may exhibit spontaneous flow reversal at the discontinuities of $\frac{dh}{dx}$.
(a) Pressure contours
(b) Velocity streamlines
Figure 5: Pressure and velocity solutions for the triangular slider with $H=0.5$. The Reynolds, ELT and PLT models overestimate the maximum velocity. Discontinuities in velocity and pressure of the ELT and $\varepsilon^{k}$-PLT models are less pronounced at smaller $H$.
(a) Pressure contours
(b) Velocity streamlines
Figure 6: Pressure and velocity solutions for the triangular slider with $H=2$. The velocity and pressure solutions of ELT and $\varepsilon^{k}$-PLT are discontinuous at the discontinuities of $\frac{dh}{dx}$. The sequence of corner eddies characteristic of the Stokes solution is not observed in any of the lubrication models.
The errors in pressure and velocity compared with the Stokes solution for the triangular slider are shown in Fig. 7. The error in velocity is very similar for all of the models. For pressure, the VA-ELT solution is a consistent improvement compared to the Reynolds and T.G.-ELT solutions, and the $\varepsilon^{k}$-PLT solutions perform better still. For negative texturing ($H<1$), the $\varepsilon^{k}$-PLT solutions of pressure have significantly smaller error relative to both the VA-ELT and Reynolds solutions; for positive texturing ($H>1$), the $\varepsilon^{k}$-PLT and VA-ELT solutions are more similar. The error in both pressure and velocity decreases as the difference between the minimum and maximum height decreases. In particular, for the case of negative texturing, the increasing magnitude of surface gradient as the minimum height decreases corresponds to an increase in error for pressure and velocity; this emphasizes the importance of considering the magnitude of surface gradients as well as the length scale ratio when selecting a fluid model.
Figure 7: Relative percent error in velocity $(u,v)$ and pressure $p(x,y)$ compared with the Stokes solution for the triangular slider with varying apex height $H$. Decreased difference between maximum and minimum height corresponds to decreased error in pressure and velocity.
5 Conclusions
We considered several models in extended lubrication theory for an incompressible, low Reynolds number fluid across a range of geometries, including several in which the standard assumptions of classical lubrication theory are not strictly satisfied. The VA-ELT model proposed here amends the velocity and pressure of the T.G.-ELT solution to ensure that incompressibility and the boundary conditions are satisfied. The solutions of velocity and pressure for the Reynolds, VA-ELT, T.G.-ELT, $\varepsilon^{k}$-PLT and Stokes models were compared for the logistic step and the triangular slider.
We found that all lubrication models have increased error in pressure and velocity as the magnitude of surface variation increases, while the length scale ration $\varepsilon=L_{x}/L_{y}$ is kept constant. In the presence of a large surface gradient, the Stokes solution exhibits significant $\frac{\partial p}{\partial y}$ in a region surrounding the surface gradient, compared with the ELT and $\varepsilon^{k}$-PLT solutions which localize this pressure variation to regions where the surface gradient is large. The consequences of this localization is overestimation of $\frac{\partial p}{\partial y}$, overestimation of the velocity magnitude, and in the case of very large surface gradients, spurious flow recirculation.
For the logistic step, the VA-ELT and $\varepsilon^{k}$-PLT solutions are an improvement on the Reynolds solution for smaller slopes; but at larger slopes, particularly for $\varepsilon^{k}$-PLT, the error exceeds that of Reynolds equation. For smaller slopes, the VA-ELT solution has the smallest error in velocity, and the $\varepsilon^{2}$-PLT solution has the smallest error in pressure. At larger slopes, the VA-ELT and $\varepsilon^{k}$-PLT solutions exhibit flow recirculation at too small of slope compared with the Stokes solution, and the velocity magnitude in the recirculation region becomes unreasonably large given the small Reynolds number.
For the triangular slider, the error in velocity is very similar for all models. For pressure, the $\varepsilon^{k}$-PLT solutions have significant improvements in error, particularly for the case of negative texturing. In the presence of the same magnitude of surface variation but at positive texturing, the VA-ELT model performs similarly to the $\varepsilon^{k}$-PLT models. This suggests that for small heights, the $\varepsilon^{k}$-PLT models are less sensitive to large or discontinuous surface gradients than the VA-ELT model. However, both the ELT and $\varepsilon^{k}$-PLT models exhibit discontinuities when the surface gradient is discontinuous. Furthermore, the sequence of corner eddies characteristic of the Stokes solution is not observed in the lubrication models.
These textured slider examples demonstrate the importance of considering both the magnitude of surface gradient and the length scale ratio when selecting a reduced order model for a low Reynolds number fluid. In the presence of only small surface gradients, significant improvements on the Reynolds solution can be achieved with an extended or perturbed model of lubrication theory; however, as demonstrated in the examples explored here, these models can be sensitive to the surface gradient.
Acknowledgments
We thank Michael Fillon for providing feedback on an early version of this manuscript. This work was supported by National Science Foundation (NSF) grant DMS-2512565 to TGF. We also acknowledge use of the Brandeis High Performance Computing Cluster (HPCC) which is partially supported by the NSF through DMR-MRSEC 2011846 and OAC-1920147.
Appendix A The Reynolds equation finite difference solution
Define the uniform discretisation of $x_{0},x_{L}]$,
(31) $\{x_{i}\}_{i=0}^{N},$ $x_{i}=x_{0}+i\Delta x,$ $\Delta x=\tfrac{1}{N}|x_{L}-x_{0}|.$
A second-order accurate difference approximation for the Reynolds equation Eq. 14 is,
(32) $\frac{1}{2\Delta x^{2}}\bigg(\Big(h_{i+1}^{3}+h_{i}^{3}\Big)p_{i+1}-\Big(h_{i+1}^{3}+2h_{i}^{3}+h_{i-1}^{3}\Big)p_{i}+\Big(h_{i}^{3}+h_{i-1}^{3}\Big)p_{i-1}\bigg)\\ =\frac{6\eta\mathcal{U}}{2\Delta x}\Big(h_{i+1}-h_{i-1}\Big).$
Appendix B The Stokes Equations
The biharmonic formulation of the incompressible Navier-Stokes equations is an effective method of solution for low Reynolds number flows [7, 9, 12, 3]. Through introduction of the stream function $\psi(x,y)$ satisfying,
(33) $u=\frac{\partial\psi}{\partial y}\hskip 30.00005ptv=-\frac{\partial\psi}{\partial x},$
the Navier-Stokes equations Eqs. 1 and 2 are expressed as,
(34) $\nabla^{4}\psi=\text{Re}\big(v\nabla^{2}u-u\nabla^{2}v).$
When $Re=0$, the velocity-stream formulation Eq. 34 reduces to the biharmonic Stokes equation $\nabla^{4}\psi=0$. The characteristic pressure used to formulate Eq. 34 is $P_{*}^{\prime}=\varepsilon P_{*}=\eta U_{*}/L_{y}$, and thus the Stokes model is not sensitive to surface variation as the lubrication models are.
To compare the Stokes solution with the one dimensional lubrication theory solution, we assume the inlet and outlet flow profiles correspond to a fully developed laminar flow with flux $\mathcal{Q}$. The surface boundary conditions for velocity are Eqs. 8 and 9, and the pressure satisfies Eq. 10. The inlet and outlet velocity profiles are,
(35) $u(x_{0},y)$ $=u_{\text{Re}}(x_{0},y)$ $\frac{\partial u}{\partial x}\Big|_{x_{L},y}=0,$
(36) $v(x_{0},y)$ $=0$ $v(x_{L},y)=0,$
where $u_{\text{Re}}(x,y)$ is the Reynolds equation velocity Eq. 11 expressed in terms of $\mathcal{Q}$ from Eq. 13. The corresponding boundary conditions for the stream function $\psi$ are,
(37) $\psi(x_{0},y)$ $=\int_{0}^{y}u_{\text{Re}}(x,\hat{y})d\hat{y}$ $\frac{\partial\psi}{\partial x}\Big|_{x_{L},y}=0,$
(38) $\psi(x,0)$ $=0$ $\psi(x,h(x))=\mathcal{Q},$
where,
(39) $\int_{0}^{y}u_{\text{Re}}(x,\hat{y})d\hat{y}=\frac{\mathcal{Q}y^{2}}{\big[h(x)\big]^{3}}\Big(3h(x)-2y\Big)+\frac{\mathcal{U}y}{\big[h(x)\big]^{2}}\Big(h(x)-y\Big)^{2}.$
The solution $\psi$, $u$, $v$ to the biharmonic Navier-Stokes equations Eqs. 34 and 33 is determined through an iterative second-order accurate finite difference method as given in [3, 7].
Once the stream and velocity have sufficiently converged, the pressure partial derivatives are determined using a centered second-order accurate finite difference discretisation of the Navier-Stokes equations Eqs. 1 and 2. For Eq. 1,
(40) $\frac{\partial p}{\partial x}\Big|_{i,j}=\frac{\nu}{\Delta x^{2}}\Big(u_{i-1,j}+u_{i+1,j}-4u_{i,j}+u_{i,j-1}+u_{i,j+1}\Big)\\ -\frac{1}{2\Delta x}\Big(u_{i,j}(u_{i+1,j}-u_{i-1,j})+v_{i,j}(u_{i,j+1}-u_{i,j-1})\Big),$
and $\frac{\partial p}{\partial y}$ as in Eq. 2 is similar. The pressure $p(x,y)$ is then determined numerically by a path integral from the outlet.
Appendix C Perturbed Lubrication Theory
The dimensionless expressions for $P_{k}$, $U_{k}$, and $V_{k}$ at orders $\varepsilon^{2}$ and $\varepsilon^{4}$ are provided. For order $\varepsilon^{2}$,
(41) $P_{2}(\bar{x},\bar{y})=-\frac{\partial U_{0}}{\partial\bar{x}}+\int_{0}^{\bar{x}}\frac{d\gamma_{3}}{d\hat{\bar{x}}}d\hat{\bar{x}}$,
(42) $U_{2}(\bar{x},\bar{y})=\frac{d^{2}\big[\bar{h}^{-3}\big]}{d\bar{x}^{2}}\Big(\bar{y}^{4}-\bar{h}^{3}\bar{y}\Big)-2\frac{d^{2}\big[\bar{h}^{-2}\big]}{d\bar{x}^{2}}\Big(\bar{y}^{3}-\bar{h}^{2}\bar{y}\Big)+\frac{1}{2}\frac{d\gamma_{3}}{d\bar{x}}(\bar{y}^{2}-\bar{h}\bar{y}),$
and,
(43) $V_{2}(\bar{x},\bar{y})=-\frac{d^{3}\big[\bar{h}^{-3}\big]}{d\bar{x}^{3}}\Big(\tfrac{1}{5}\bar{y}^{5}-\tfrac{1}{2}\bar{h}^{3}\bar{y}^{2}\Big)+\frac{3}{2}\frac{d^{2}\big[\bar{h}^{-3}\big]}{d\bar{x}^{2}}\frac{d\bar{h}}{d\bar{x}}\bar{h}^{2}\bar{y}^{2}\\ +\frac{d^{3}\big[\bar{h}^{-2}\big]}{d\bar{x}^{3}}\Big(\tfrac{1}{2}\bar{y}^{4}-\bar{h}^{2}\bar{y}^{2}\Big)-2\frac{d^{2}\big[\bar{h}^{-2}\big]}{d\bar{x}^{2}}\frac{d\bar{h}}{d\bar{x}}\bar{h}\bar{y}^{2}\\ -\frac{1}{2}\frac{d^{2}\gamma_{3}}{d\bar{x}^{2}}\Big(\tfrac{1}{3}\bar{y}^{3}-\tfrac{1}{2}\bar{h}\bar{y}^{2}\Big)+\frac{1}{4}\frac{d\gamma_{3}}{d\bar{x}}\frac{d\bar{h}}{d\bar{x}}\bar{y}^{2},$
where,
(44) $\frac{d\gamma_{3}}{d\bar{x}}=-\frac{18}{5}\frac{d^{2}\big[\bar{h}^{-3}\big]}{d\bar{x}^{2}}\bar{h}^{2}+6\frac{d^{2}\big[\bar{h}^{-2}\big]}{d\bar{x}^{2}}\bar{h}.$
And likewise for order $\varepsilon^{4}$,
(45) $P_{4}(\bar{x},\bar{y})=-\frac{\partial U_{2}}{\partial\bar{x}}+\int_{0}^{\bar{y}}\frac{\partial^{2}V_{0}}{\partial{\bar{x}^{2}}}d\hat{\bar{y}}+\int_{0}^{\bar{x}}\frac{d\gamma_{5}}{d\hat{\bar{x}}}d\hat{\bar{x}}$,
(46) $U_{4}(\bar{x},\bar{y})=-\frac{1}{20}\frac{d^{4}\big[\bar{h}^{-3}\big]}{d\bar{x}^{4}}\Big(\bar{y}^{6}-\bar{h}^{5}\bar{y}\Big)+\frac{3}{20}\frac{d^{4}\big[\bar{h}^{-2}]}{d\bar{x}^{4}}\Big(\bar{y}^{5}-\bar{h}^{4}\bar{y}\Big)\\ +\Bigg(\frac{1}{3}\frac{d^{2}\phi_{1}}{d\bar{x}^{2}}-\frac{2}{3}\frac{d^{2}\phi_{2}}{d\bar{x}^{2}}+\frac{1}{6}\frac{d^{2}\phi_{3}}{d\bar{x}^{2}}\Bigg)\Big(\bar{y}^{3}-\bar{h}^{2}\bar{y}\Big)\\ -\frac{1}{12}\frac{d^{3}\gamma_{3}}{d\bar{x}^{3}}\Big(\bar{y}^{4}-\bar{h}^{3}\bar{y}\Big)+\frac{1}{2}\frac{d\gamma_{5}}{d\bar{x}}\Big(\bar{y}^{2}-\bar{h}\bar{y}\Big),$
and,
(47) $V_{4}(\bar{x},\bar{y})=-\frac{1}{20}\frac{d^{5}\big[\bar{h}^{-3}\big]}{d\bar{x}^{5}}\Big(\tfrac{1}{7}\bar{y}^{7}-\tfrac{1}{2}\bar{h}^{5}\bar{y}^{2}\Big)+\frac{1}{8}\frac{d^{4}\big[\bar{h}^{-3}\big]}{d\bar{x}^{4}}\frac{d\bar{h}}{d\bar{x}}\bar{h}^{4}\bar{y}^{2}\\ +\frac{3}{20}\frac{d^{5}\big[\bar{h}^{-2}\big]}{d\bar{x}^{5}}\Big(\tfrac{1}{6}\bar{y}^{6}-\tfrac{1}{2}\bar{h}^{4}\bar{y}^{2}\Big)-\frac{3}{10}\frac{d^{4}\big[\bar{h}^{-2}\big]}{d\bar{x}^{4}}\frac{d\bar{h}}{d\bar{x}}\bar{h}^{3}\bar{y}^{2}\\ +\frac{1}{3}\Bigg(\frac{d^{3}\phi_{1}}{d\bar{x}^{3}}-2\frac{d^{3}\phi_{2}}{d\bar{x}^{3}}\Bigg)\Big(\tfrac{1}{4}\bar{y}^{4}-\tfrac{1}{2}\bar{h}^{2}\bar{y}^{2}\Big)-\Bigg(\frac{d^{2}\phi_{1}}{d\bar{x}^{2}}-2\frac{d^{2}\phi_{2}}{d\bar{x}^{2}}\Bigg)\frac{d\bar{h}}{d\bar{x}}\bar{h}\bar{y}^{2}\\ +\frac{1}{6}\Bigg(\frac{d^{3}\phi_{3}}{d\bar{x}^{3}}\Big(\tfrac{1}{4}\bar{y}^{4}-\tfrac{1}{2}\bar{h}^{2}\bar{y}^{2}\Big)-\frac{d^{2}\phi_{3}}{d\bar{x}^{2}}\frac{d\bar{h}}{d\bar{x}}\bar{h}\bar{y}^{2}\Bigg)-\frac{1}{12}\frac{d^{4}\gamma_{3}}{d\bar{x}^{4}}\Big(\tfrac{1}{5}\bar{y}^{5}-\tfrac{1}{2}\bar{h}^{3}\bar{y}^{2}\Big)\\ +\frac{1}{8}\frac{d^{3}\gamma_{3}}{d\bar{x}^{3}}\frac{d\bar{h}}{d\bar{x}}\bar{h}^{2}\bar{y}^{2}+\frac{1}{2}\frac{d^{2}\gamma_{5}}{d\bar{x}^{2}}\Big(\tfrac{1}{3}\bar{y}^{3}-\tfrac{1}{2}\bar{h}\bar{y}^{2}\Big)-\frac{1}{4}\frac{d\gamma_{5}}{d\bar{x}}\frac{d\bar{h}}{d\bar{x}}\bar{y}^{2},$
where,
(48) $\frac{d^{2}\phi_{1}}{d\bar{x}^{2}}=\frac{d^{4}\big[\bar{h}^{-3}]}{d\bar{x}^{4}}\bar{h}^{3}+6\frac{d^{3}\big[\bar{h}^{-3}\big]}{d\bar{x}^{3}}\frac{d\bar{h}}{d\bar{x}}\bar{h}^{2}+\frac{d^{2}\big[\bar{h}^{-3}\big]}{d\bar{x}^{2}}\Bigg(6\bigg[\frac{d\bar{h}}{d\bar{x}}\bigg]^{2}\bar{h}+3\frac{d^{2}\bar{h}}{d\bar{x}^{2}}\bar{h}^{2}\Bigg),$
(49) $\frac{d^{2}\phi_{2}}{d\bar{x}^{2}}=\frac{d^{4}\big[h^{-2}\big]}{d\bar{x}^{4}}\bar{h}^{2}+4\frac{d^{3}\big[\bar{h}^{-2}\big]}{d\bar{x}^{3}}\frac{d\bar{h}}{d\bar{x}}\bar{h}+2\frac{d^{2}\big[\bar{h}^{-2}\big]}{d\bar{x}^{2}}\Bigg(\bigg[\frac{d\bar{h}}{d\bar{x}}\bigg]^{2}+\frac{d^{2}\bar{h}}{d\bar{x}^{2}}\bar{h}\Bigg),$
(50) $\frac{d^{2}\phi_{3}}{d\bar{x}^{2}}=\frac{d^{3}\gamma_{3}}{d\bar{x}^{3}}\bar{h}+2\frac{d^{2}\gamma_{3}}{d\bar{x}^{2}}\frac{d\bar{h}}{d\bar{x}}+\frac{d^{2}\bar{h}}{d\bar{x}^{2}}\frac{d\gamma_{3}}{d\bar{x}},$
and,
(51) $\frac{d\gamma_{5}}{d\bar{x}}=\frac{3}{14}\frac{d^{4}\big[\bar{h}^{-3}\big]}{d\bar{x}^{4}}\bar{h}^{4}-\frac{3}{5}\frac{d^{4}\big[\bar{h}^{-2}]}{d\bar{x}^{4}}\bar{h}^{3}+\frac{3}{10}\frac{d\gamma_{3}}{dx}\bar{h}^{2}.$
Appendix D The Backward Facing Step
In the limit $\lambda\to\infty$, the logistic step Eq. 29 resembles the classical backward facing step (BFS) centered at $L/2$,
(52) $h(x)=\begin{cases}H_{\text{in}}&\hskip 20.00003pt0\leq x\leq L/2\\ H_{\text{out}}&\hskip 20.00003ptL/2<x\leq L\end{cases}.$
Due to the BFS having piecewise zero surface gradient, the solutions according to the Reynolds equation, ELT and $\varepsilon^{k}$-PLT models are all equivalent. The pressure contours of this solution are purely one dimensional, and the velocity streamlines are everywhere parallel to $y=0$ instead of wrapping around the sharp corners of the step. In contrast, the Stokes solution is able to capture the discontinuity in the height, leading to a similar flow profile as in the logistic step at large $\lambda$.
The Stokes solutions of pressure and velocity for the BFS and the logistic step at $\lambda=32$ are shown in Fig. 8. The two configurations have the same $H_{\text{in}}/H_{\text{out}}=2$, $L=16$, $\nu=1$, $\mathcal{Q}=1$ and $\mathcal{U}=0$. According to the Stokes solution, the BFS and the logistic step at large $\lambda$ have similar total pressure drops and flow structures. In particular, flow recirculation is significant in the logistic step for $\lambda>16$, with similar positioning to the recirculation observed in the BFS. Yet, as seen in Fig. 2, the Reynolds solution does not exhibit flow recirculation in the logistic step for any $\lambda$, and the VA-ELT and $\varepsilon^{k}$-PLT solutions exhibit flow recirculation at too small of $\lambda$ and with too fast of velocities compared with the Stokes solution.
(a) Pressure contours $p(x,y)$
(b) Velocity streamlines $(u,v)$
Figure 8: The Stokes solutions to the logistic step with $\lambda=32$ (left) and the BFS (right). The logistic step at large $\lambda$ exhibits similar patterns of corner flow recirculation and $\frac{\partial p}{\partial y}$ as in the BFS.